Find equation of the line which is perpendicular to the line
y=5x+9 and has y-intercept = 6.

Answer by mananth(16946)

y = 5 x + 9
Compare this equation with y=mx+b
slope m = 5.00

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be -0.20=> -1/5

m= - 1/5 ,point ( 0 , 6 )
Find b by plugging the values of m & the point in
y=mx+b
6 = + b
b= 6
m= - 1/5
The required equation is y =- 1/ 5x+ 6
m.ananth@hotmail.ca